OpenStudy (olsolneo123):

The function h(x) = 106x2 + 110x + 495 can also be written as which of the following?

OpenStudy (olsolneo123):

A. h(x) + 495 = 106x2 + 110x B. y = 106x2 + 110x - 495 C. y + 495 = 106x2 + 110x D. y = 106x2 + 110x + 495

OpenStudy (anonymous):

D.

OpenStudy (olsolneo123):

thank you!:)

OpenStudy (anonymous):

You are welcome.